Zeus Corporation produces cultured diamonds via a secretive process that grows the diamonds in a vacuum chamber filled with a carbon gas cloud. The diamonds are produced in a single continuous process and Zeus uses the weighted-average process costing method of accounting for production.

Below is the company's calculation of cost per equivalent unit for October. During October, the company

completed and transferred 8,000 diamonds to finished goods. An additional 4,000 units were still in process at the end of the month. The ending work in process was 60% complete with respect to direct materials and 40% complete with respect to both elements of conversion cost.

Prepare a schedule showing the allocation of total cost between finished goods and ending work in process.

Costs per equivalent unit :

Conversion

Total Cost

Direct Materials

Direct Labor

Factory Overhead

Beginning Work in Process

$3,900,000

$1,170,000

$780,000

$1,950,000

Cost incurred during period

9,300,000

2,000,000

2,325,000

$5,115,000

Total cost

$13,340,000

$3,170,000

$3,105,000

$7,065,000

Equivalent units

10,400

9,600

9,600

Costs per equivalent unit

$304.81

$323.44

$735.94

$1,059.38

$1,364.19

A.

Transferred to Finished Goods

$12,831,980

Total Ending Work in Process

$3,029,997

Total Cost Allocation

$15,861,894

B.

Transferred to Finished Goods

$8,831,354

Total Ending Work in Process

$5,029,231

Total Cost Allocation

$13,860,585

C.

Transferred to Finished Goods

$10,913,520

Total Ending Work in Process

$2,426,552

Total Cost Allocation

$13,340,072

D.

Transferred to Finished Goods

$14,833,934

Total Ending Work in Process

$1,229,566

Total Cost Allocation

$16,863,500

E.

Transferred to Finished Goods

$12,831,574

Total Ending Work in Process

$8,029,903

Total Cost Allocation

$18,861,875
Answer #1
materials CC
completed & transferred 8000 8000
Ending work in process 2400 1600
Equivalent units 10400 9600
Transferred to finished goods 8000*1364.183= 10913520
ending work in process
Materials 2400*304.81= 731544
direct labor 1600*323.44= 517504
conversion 1600*735.9375 1177504 2426552
total cost allocation 13340072
answer) option c